# CustomVideo: Customizing Text-to-Video Generation with Multiple Subjects

### v2

@@ Abstract
Customized text-to-video generation aims to generate high-quality videos guided by text prompts and subject references. Current approaches for personalizing text-to-video generation suffer from tackling multiple subjects, which is a more challenging and practical scenario. In this work, our aim is to promote multi-subject guided text-to-video customization. We propose CustomVideo, a novel framework that can generate identity-preserving videos with the guidance of multiple subjects. To be specific, firstly, we encourage the co-occurrence of multiple subjects via composing them in a single image. Further, upon a basic text-to-video diffusion model, we design a simple yet effective attention control strategy to disentangle different subjects in the latent space of diffusion model. Moreover, to help the model focus on the specific area of the object, we segment the object from given reference images and provide a corresponding object mask for attention learning. Also, we collect a multi-subject text-to-video generation dataset as a comprehensive benchmark, with 63 individual subjects from 13 different categories and 68 meaningful pairs. Extensive qualitative, quantitative, and user study results demonstrate the superiority of our method compared to previous state-of-the-art approaches. The project page is https://kyfafyd.wang/projects/customvideo.
1 Introduction
@@
A recent work, VideoDreamer [12], proposes disen-mix finetuning and human-in-the-loop re-finetuning strategies based on Stable Diffusion [13], aiming to generate videos from multiple subjects. However, VideoDreamer falls short in guaranteeing the co-occurrence of multiple subjects and disentangling similar subjects due to its inconsistent object mixing strategy. As shown in Figure 3, both the car and cat can not be consistently generated across all frames. Another straight forward solution for multi-subject driven T2V generation is by combining multi-subject driven Text-to-Image (T2I) and Image-to-Video (I2V) models. However, this naive approach fails from two aspects, including inaccurate T2I customization and lack of motion (see Figure 3 ‘DisenDiff [14] + SVD [15]’).
In contrast, in our approach, we ensure the co-occurrence of multiple objects during model training, which encourages the model to capture the presence of different subjects simultaneously, thereby facilitating co-occurrence during inference. Additionally, we propose an attention control mechanism to disentangle multiple subjects during training, effectively guiding the model to focus on the corresponding subject area while disregarding irrelevant parts of the image. To facilitate this process, we incorporate a ground truth object mask, obtained through segmentation either from a model like SAM [16] or provided by human annotators, as supervision during optimization. Our attention mechanism consists of two key designs that contribute to the disentanglement of subjects. Firstly, we highlight the corresponding subject area using the ground truth object mask on the cross-attention map, promoting better preservation of subject identity. Secondly, we optimize the cross-attention map towards a slight negative value, excluding the desired subject and mitigating the influence of irrelevant area in the input image. To comprehensively evaluate our proposed approach, we have curated a diverse dataset covering a wide range of 13 categories. Beyond the MultiStudioBench [12] which includes 12 subject pairs, our dataset comprises 63 individual subjects and 68 subject pairs, notably featuring some challenging scenarios involving visually similar objects. Through extensive experiments on this benchmark dataset, we provide qualitative, quantitative, and user study results that demonstrate the superiority of our method in generating high-quality videos with customized subjects. In summary, our contributions are as follows:
@@
Our method consistently outperforms previous state-of-the-art (SOTA) approaches in qualitative, quantitative and human preference evaluation, e.g., 11.99% CLIP Image Aligenment and 23.39% DINO Image Aligenment better than the previous best method.
2 Related Work
Text-to-Video Generation.
@@ 2 Related Work
Text-to-video generation has made significant advancements in recent years [17, 18, 1]. Early approaches [19, 20, 21, 22] employed GANs [23] and VQ-VAE [24], while more recent works have explored diffusion models to generate high-quality videos [25, 26, 27, 28]. Make-A-Video [29] utilizes a pre-trained image diffusion model with additional temporal attention finetuning. VideoLDM [30] introduces a multi-stage alignment approach in the latent space to generate high-resolution and temporally consistent videos. Other methods [31, 32] generate videos with an image as the first frame and randomly initialized subsequent frames. To enhance controllability, VideoComposer [33] incorporates additional guidance signals, such as depth maps, to produce desired videos alongside text inputs. VideoDirectorGPT [34] aims to control the video generation with the guidance from different scenes and specific layouts generated by GPT4 in the temporal axis. GEST [35] models the representation between the text and video via a graph of events in the spatial and temporal space, in which the video is generated following a timeline. Tune-A-Video [36] proposes a temporal self-attention module that fine-tunes a pre-trained image diffusion model, achieving successful generation of videos with specific text guidance. Furthermore, diffusion-based video-to-video editing approaches [37, 38] have also been proposed for practical usages.
Subject-driven Customization.
@@
There has been a growing interest in customizing pre-trained image and video diffusion models for personalized generation [39, 40, 41, 42, 43, 44, 11]. Customization involves generating images and videos with specific subjects, typically based on a few reference images. For image diffusion customization, Textual Inversion [39] represents a specific object as a learnable text token using only a few reference images. This learned text token can then be integrated into a sentence to generate personalized images during inference. Additionally, DreamBooth [45] fine-tunes the weights of the diffusion model to improve fidelity in image generation. Several works [46, 14, 47, 48] have explored personalized image diffusion with multiple subjects, focusing on parameter-efficient finetuning and text embedding learning. While there have been initial attempts to customize video diffusion, such as VideoAssembler [9], VideoBooth [10], and ID-Animator [49], which use reference images to personalize the video diffusion model while preserving subject identity, and DreamVideo [11], which decouples the learning process for subject and motion customization, these methods are designed for single objects and cannot handle multiple subjects when given. A recent work, VideoDreamer [12], proposes multi-subject driven video customization through disen-mix finetuning strategy with LoRA [50]. However, the generated videos do not guarantee the co-occurrence of multiple subjects, or disentanglement of different subjects. In this work, we propose a simple yet effective co-occurrence and attention mechanism that disentangles multiple subjects using masks as guidance while preserving the co-occurrence of subjects in the generated videos.
3 Method
3.1 Preliminary: Text-to-Video
@@
3.2 CustomVideo with Multiple Subjects
@@
An overview of our proposed CustomVideo framework is illustrated in Figure 2. For customization perspective, we introduce a new learnable word token for every subject, e.g., ‘<new1>’ for cat and ‘<new2>’ for dog, representing the corresponding identity. During the training stage, we employ a concatenation technique to ensure the co-occurrence of multiple subjects. Specifically, we combine these subjects into a single image, facilitating model’s learning of multi-subject patterns. To address the challenge of entanglement among highly similar subjects, we propose an attention control mechanism. This mechanism ensures that the learnable word tokens align with their corresponding regions on the cross-attention map (see Figure 6). By achieving this alignment, we enable the model to disentangle different subjects and enhance the quality of the generated videos. The training process focuses on training the subject-related learnable word tokens, as well as the key and value weights within the cross-attention layers of the U-Net architecture, adopting a parameter-efficient finetuning approach. During the inference stage, users only need to provide a text prompt description integrated with the corresponding learned word tokens to generate high-quality videos aligned with their preferences. In the following, we will delve into the detailed workings of CustomVideo.
Co-occurrence Control.
@@
In the context of multi-subject driven T2V generation, ensuring that the model consistently learns to generate videos with multiple subjects is crucial. In a previous work, VideoDreamer [12], the authors proposed a disen-mix strategy for multi-subject driven generation by fine-tuning the model using both single subject images and concatenated images of multiple subjects. However, we have identified that this mixing strategy can confuse the model due to the inconsistent number of subjects present in a single image. Consequently, it leads to unstable generation of videos with multiple subjects (see line ‘both single and concat’ in Figure 5). In our approach, we have found that fine-tuning the model using only concatenated images of multiple subjects is sufficient to ensure the co-occurrence of multiple subjects in the generated videos. Additionally, we have observed that providing clear subjects without background aids the model in learning the specific characteristics of the subjects more effectively. To achieve this, we perform background removal on the subject images. This can be accomplished manually or by employing an automatic tool such as the SAM model [16].
Attention Control.
@@
Ensuring the co-occurrence of multiple subjects in generated videos is achieved through the above. However, a more challenging task is preserving the distinct characteristics of each subject without interference. Simply fine-tuning the model with concatenated images can lead to confusion among subject characteristics. As shown in Figure 5 (line ‘w/o pos. attn.’), the generated cat predominantly resembles the shape of the provided dog sample, despite having similar color and texture to the provided cat. Thus, disentangling the multiple subjects becomes crucial to generate high-quality videos that faithfully represent each subject’s characteristics.
@@
In our approach, the learnable word tokens operate in the cross-attention layers of the diffusion model. To effectively regulate the subject learning process, we can directly leverage the cross-attention map. As illustrated in Figure 2, we employ an automatic segmentor to obtain the ground truth mask for each subject, which indicates the spatial position of the subjects. During the training stage, we extract the cross-attention map from the cross-attention layer using the activations of each word in the given text prompt. For each learnable word token, such as ‘new1’, we enhance the corresponding area on the cross-attention map by encouraging the alignment between the subject area and the ground truth subject mask with a loss function as the following:
@@
where is the number of subjects, the corresponding area of subject in the mask is filled by value 1 while the remaining area is filled by 0. By employing this positive attention mechanism, the model is compelled to allocate more attention to the correct subject area, leading to an effective learning of the corresponding subject characteristics.
@@
The positive-style attention mechanism mentioned above effectively enhances the learning of specific characteristics for each subject. However, there exist issues with the generated subjects, particularly when irrelevant area is present. For example, in Figure 5 (line ‘w/o neg. attn.’), we can observe that the generated dog’s legs are affected by color information from the given cat, which is not desirable. To address this problem, we introduce negative guidance by considering the area outside the subject. In addition to the positive guidance within the ground truth subject mask, we incorporate a small negative value, denoted as , into the region outside the subject within the mask . This modified mask, denoted as , is then used in Eq. (2) to regulate the subject learning process. By integrating negative guidance, we can alleviate the issue of irrelevant area and improve the fidelity of the generated subjects.
3.3 Model Training and Inference
Training Strategy.
@@ 3.3 Model Training and Inference
During training, we only fine-tune the weights of key and value in all of the cross-attention layers. We use a more flexible approach by fine-tuning the model with subject images extended to single-frame videos as both input and supervision, rather than using subject videos. This approach guarantees the co-occurrence of multiple subjects because subject images are much easier to obtain and smoother to concatenate compared to subject videos. Moreover, fine-tuning with subject images saves lots of computation cost. Inspired by previous T2I personalization [45, 40], we conduct class-specific prior preservation to improve the diversity of generated videos and alleviate the issue of language drift. The loss for prior preservation is
@@
where is the latent code of the input class image, is the noise prediction from the model , is the randomly sampled Gaussian noise, is the text condition for class image, and is the sampling timestep. To this end, we train our CustomVideo via an end-to-end manner, with the following overall training objective:
where and are two hyper-parameters to control the weight of attention control and prior preservation, respectively.
Inference.

### v3

@@ Abstract
Customized text-to-video generation aims to generate high-quality videos guided by text prompts and subject references. Current approaches for personalizing text-to-video generation suffer from tackling multiple subjects, which is a more challenging and practical scenario. In this work, our aim is to promote multi-subject guided text-to-video customization. We propose CustomVideo, a novel framework that can generate identity-preserving videos with the guidance of multiple subjects. To be specific, firstly, we encourage the co-occurrence of multiple subjects via composing them in a single image. Further, upon a basic text-to-video diffusion model, we design a simple yet effective attention control strategy to disentangle different subjects in the latent space of diffusion model. Moreover, to help the model focus on the specific area of the object, we segment the object from given reference images and provide a corresponding object mask for attention learning. Also, we collect a multi-subject text-to-video generation dataset as a comprehensive benchmark. Extensive qualitative, quantitative, and user study results demonstrate the superiority of our method compared to previous state-of-the-art approaches. The project page is https://kyfafyd.wang/projects/customvideo.
I Introduction
@@
A recent work, VideoDreamer [12], proposes disen-mix finetuning and human-in-the-loop re-finetuning strategies based on Stable Diffusion [13], aiming to generate videos from multiple subjects. However, VideoDreamer falls short in guaranteeing the co-occurrence of multiple subjects and disentangling similar subjects due to its inconsistent object mixing strategy. As shown in Fig.4, both the car and cat can not be consistently generated across all frames. Another straight forward solution for multi-subject driven T2V generation is by combining multi-subject driven Text-to-Image (T2I) and Image-to-Video (I2V) models. However, this naive approach fails from two aspects, including inaccurate T2I customization and lack of motion (see Fig.4 ‘DisenDiff [14] + SVD [3]’).
In contrast, in our approach, we ensure the co-occurrence of multiple objects during model training for the first time in multi-subject driven T2V customization, which encourages the model to capture the presence of different subjects simultaneously, thereby facilitating co-occurrence during inference. Additionally, we propose an attention control mechanism to disentangle multiple subjects during training, effectively guiding the model to focus on the corresponding subject area while disregarding irrelevant parts of the image. To facilitate this process, we incorporate a ground truth object mask, obtained through segmentation either from a model like SAM [15] or provided by human annotators, as supervision during optimization. Our attention mechanism consists of two key designs that contribute to the disentanglement of subjects. Firstly, we highlight the corresponding subject area using the ground truth object mask on the cross-attention map, promoting better preservation of subject identity. Secondly, we optimize the cross-attention map towards a slight negative value, excluding the desired subject and mitigating the influence of irrelevant area in the input image. To comprehensively evaluate our proposed approach, we have curated a diverse dataset covering a wide range of categories. Through extensive experiments on this benchmark dataset, we provide qualitative, quantitative, and user study results that demonstrate the superiority of our method in generating high-quality videos with customized subjects. In summary, our contributions are as follows:
@@
Our method consistently outperforms previous state-of-the-art (SOTA) approaches in qualitative, quantitative and human preference evaluation, e.g., 11.99% CLIP Image Alignment and 23.39% DINO Image Alignment better than the previous best method.
II Related Work
Text-to-Video Generation
@@ II Related Work
Text-to-video generation has made significant advancements in recent years [16, 17, 1]. Early approaches [18, 19, 20, 21, 22] employed GANs [23] and VQ-VAE [24], while more recent works have explored diffusion models to generate high-quality videos [25, 26, 27, 28, 29, 30, 31]. Make-A-Video [32] utilizes a pre-trained image diffusion model with additional temporal attention finetuning. VideoLDM [33] introduces a multi-stage alignment approach in the latent space to generate high-resolution and temporally consistent videos. Other methods [34, 35] generate videos with an image as the first frame and randomly initialized subsequent frames. To enhance controllability, VideoComposer [36] incorporates additional guidance signals, such as depth maps, to produce desired videos alongside text inputs. VideoDirectorGPT [37] aims to control the video generation with the guidance from different scenes and specific layouts generated by GPT4 in the temporal axis. GEST [38] models the representation between the text and video via a graph of events in the spatial and temporal space, in which the video is generated following a timeline. Tune-A-Video [39] proposes a temporal self-attention module that fine-tunes a pre-trained image diffusion model, achieving successful generation of videos with specific text guidance. Furthermore, diffusion-based video-to-video editing approaches [40, 41] have also been proposed for practical usages.
Subject-driven Customization
@@
There has been a growing interest in customizing pre-trained image and video diffusion models for personalized generation [42, 43, 44, 45, 46, 47, 48, 11]. Customization involves generating images and videos with specific subjects, typically based on a few reference images. For image diffusion customization, Textual Inversion [42] represents a specific object as a learnable text token using only a few reference images. This learned text token can then be integrated into a sentence to generate personalized images during inference. Additionally, DreamBooth [49] fine-tunes the weights of the diffusion model to improve fidelity in image generation. Several works [50, 14, 51, 52] have explored personalized image diffusion with multiple subjects, focusing on parameter-efficient finetuning and text embedding learning. While there have been initial attempts to customize video diffusion, such as VideoAssembler [9], VideoBooth [10], and ID-Animator [53], which use reference images to personalize the video diffusion model while preserving subject identity, and DreamVideo [11], which decouples the learning process for subject and motion customization, these methods are designed for single objects and cannot handle multiple subjects when given. A recent work, VideoDreamer [12], proposes multi-subject driven video customization through disen-mix finetuning strategy with LoRA [54]. However, the generated videos do not guarantee the co-occurrence of multiple subjects, or disentanglement of different subjects. An alternative way is to combine multi-subject T2I customization and I2V models, such as DisenDiff [14] + SVD-XT [3]. However, such approach suffers from two fundamental limitations: 1) lack of spatial co-occurrence training: T2I models are typically trained on single-image supervision, lacking explicit signals to enforce the spatial co-occurrence of multiple subjects. In contrast, our CustomVideo trains on concatenated multi-subject images, forcing the model to learn their joint spatial distributions from the outset; 2) motion semantic gap: I2V models learn motion patterns from generic video data but lack subject-specific motion priors. For instance, in Figure 4, “DisenDiff+SVD-XT” generates static or disjointed motions for the car and barn, failing to capture dynamic interactions like “stopping”. In this work, we propose a simple yet effective co-occurrence and attention mechanism that disentangles multiple subjects using masks as guidance while preserving the co-occurrence of subjects in the generated videos.
III Method
III-A Preliminary: Text-to-Video
@@
III-B CustomVideo with Multiple Subjects
@@
An overview of our proposed CustomVideo framework is illustrated in Fig. 2. For customization perspective, we introduce a new learnable word token for every subject, e.g., ‘new1’ for cat and ‘new2’ for dog, representing the corresponding identity. During the training stage, we employ a concatenation technique to ensure the co-occurrence of multiple subjects. Specifically, we combine these subjects into a single image, facilitating model’s learning of multi-subject patterns. To address the challenge of entanglement among highly similar subjects, we propose an attention control mechanism. This mechanism ensures that the learnable word tokens align with their corresponding regions on the cross-attention map (see Fig. 9). By achieving this alignment, we enable the model to disentangle different subjects and enhance the quality of the generated videos. The training process focuses on training the subject-related learnable word tokens, as well as the key and value weights within the cross-attention layers of the U-Net architecture, adopting a parameter-efficient finetuning approach. During the inference stage, users only need to provide a text prompt description integrated with the corresponding learned word tokens to generate high-quality videos aligned with their preferences. In the following, we will delve into the detailed workings of CustomVideo.
Co-occurrence Control
@@
In the context of multi-subject driven T2V generation, ensuring that the model consistently learns to generate videos with multiple subjects is crucial. In a previous work, VideoDreamer [12], the authors proposed a disen-mix strategy for multi-subject driven generation by fine-tuning the model using both single subject images and concatenated images of multiple subjects. However, we have identified that this mixing strategy can confuse the model due to the inconsistent number of subjects present in a single image. Consequently, it leads to unstable generation of videos with multiple subjects (see line ‘both single and concat’ in Fig. 10). In our approach, we have found that fine-tuning the model using only concatenated images of multiple subjects is sufficient to ensure the co-occurrence of multiple subjects in the generated videos. Additionally, we have observed that providing clear subjects without background aids the model in learning the specific characteristics of the subjects more effectively. To achieve this, we perform background removal on the subject images. This can be accomplished manually or by employing an automatic tool such as the SAM model [15].
Attention Control
@@
Ensuring the co-occurrence of multiple subjects in generated videos is achieved through the above. However, a more challenging task is preserving the distinct characteristics of each subject without interference. Simply fine-tuning the model with concatenated images can lead to confusion among subject characteristics. As shown in Fig. 10 (line ‘w/o pos. attn.’), the generated cat predominantly resembles the shape of the provided dog sample, despite having similar color and texture to the provided cat. Thus, disentangling the multiple subjects becomes crucial to generate high-quality videos that faithfully represent each subject’s characteristics.
@@
In our approach, the learnable word tokens operate in the cross-attention layers of the diffusion model. To effectively regulate the subject learning process, we can directly leverage the cross-attention map. As illustrated in Fig. 2, we employ an automatic segmentor to obtain the ground truth mask for each subject, which indicates the spatial position of the subjects. During the training stage, we extract the cross-attention map from the cross-attention layer using the activations of each word in the given text prompt. For each learnable word token, such as ‘new1’, we enhance the corresponding area on the cross-attention map by encouraging the alignment between the subject area and the ground truth subject mask with a loss function as the following:
@@
where is the number of subjects, is the training batch size, and the corresponding area of subject in the mask is filled by value 1 while the remaining area is filled by 0. By employing this positive attention mechanism, the model is compelled to allocate more attention to the correct subject area, leading to an effective learning of the corresponding subject characteristics.
@@
The positive-style attention mechanism mentioned above effectively enhances the learning of specific characteristics for each subject. However, there exist issues with the generated subjects, particularly when irrelevant area is present. For example, in Fig. 10 (line ‘w/o neg. attn.’), we can observe that the generated dog’s legs are affected by color information from the given cat, which is not desirable. To address this problem, we introduce negative guidance by considering the area outside the subject. In addition to the positive guidance within the ground truth subject mask, we incorporate a small negative value, denoted as , into the region outside the subject within the mask . This modified mask, denoted as , is then used to replace in the attention loss. So Eq. (2) becomes
By integrating negative guidance, we can alleviate the issue of irrelevant area and improve the fidelity of the generated subjects.
III-C Model Training and Inference
Training Strategy
@@ III-C Model Training and Inference
During training, we only fine-tune the weights of key and value in all of the cross-attention layers. We use a more flexible approach by fine-tuning the model with subject images extended to single-frame videos as both input and supervision, rather than using subject videos. This approach guarantees the co-occurrence of multiple subjects because subject images are much easier to obtain and smoother to concatenate compared to subject videos. Moreover, fine-tuning with subject images saves lots of computation cost. Inspired by previous T2I personalization [49, 43], we conduct class-specific prior preservation to improve the diversity of generated videos and alleviate the issue of language drift. Specifically, we generate 200 class images with Stable Diffusion v2.1 [13] for each subject, in which the generating prompts are obtained from Claude-3-Opus [56]. Then, to guarantee the co-occurrence of multiple subjects, similarly, we concatenate the class images of multiple subjects. The concatenated class image is then fed into the U-Net for training with the text prompt “a cat and a dog” (taking cat and dog as an example). To be note that these images are also conducted background removal. As there is no learnable text tokens during prior preservation training, the attention loss is not applied. We only apply the reconstruction loss for supervision as the following:
@@
where is the latent code of the input concatenated class image, is the noise prediction from the model , is the randomly sampled Gaussian noise, is the text condition for the concatenated class image, and is the sampling timestep. To this end, we train our CustomVideo via an end-to-end manner, with the following overall training objective:
where and are two hyper-parameters to control the weight of attention control and prior preservation, respectively. Our proposed co-occurrence and attention control mechanism is exclusively applied during model training to assist the model and learnable text tokens in disentangling the identities of multiple subjects. During the training stage, images of different subjects are carefully concatenated without overlap, enabling the attention control mechanism to accurately locate the regions of each subject and learn their corresponding features.
Inference